Write the equation of the line with slope 5 and y intercept (0,-3)
Answer by rapaljer(4551)

The equation of a straight line can always be written in the form of y = mx +b. That is, there is a number times x, then plus (or minus) some other number. The number that is times x is called the "coefficient" of x, and this number turns out to be the slope of the line, and we usually call this number "m". The other number, the one that doesn't have the x in it, is called the "number term", and this number turns out to be the y-intercept, which is the point where the graph crosses the y-axis. The y-intercept is also the value of y that you get when you let x=0. We call the y-intercept b. Put it together, and that's why we write the equation y=mx+b.

So, in your problem, you were given the slope and the y-intercept. Your problem was "write the equation of the line with slope 5 and y intercept (0,-3)." Now, in the formula "y=mx+b" if the slope is 5, then m=5. and if the y-intercept is (0,-3), then b=-3, so y =mx+b means that
y=5x+(-3) or y=5x-3.

See if these details help you understand the problem.

R^2 at SCC